OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation x + 2y = 4. x-intercept: (0, 2), y-intercept: (4, 0) x-intercept: (2, 0), y-intercept: (4, 0) x-intercept: (0, 4), y-intercept: (2, 0) x-intercept: (4, 0), y-intercept: (0, 2)

OpenStudy (anonymous):

To find the x-intercept, let y = 0: x + 2(0) = 4 x + 0 = 4 Therefore, x = 4 The coordinate is (4,0) To find the y-intercept, let x=0. 0 + 2y = 4 2y = 4 Divide both sides by 2. Therefore, 4 divided by 2 is 2, and y=2. The coordinate is (0,2) Therefore D (last choice) is your answer.